PLEASE HELP. I NEED HELP IN STATISTICS. THANK YOU SO MUCH
The Web-based company gift baskets galore! Gifts have a goal of processing 95 percent of its orders on the same day they are received. If 570 out of the next 600 orders are processed on the same day, would this prove that they are exceeding their goal, using α = .05? Explain your decision. clearly state your decision on the null hypothesis.

Answer by stanbon(75887)

The Web-based company gift baskets galore! Gifts have a goal of processing 95 percent of its orders on the same day they are received. If 570 out of the next 600 orders are processed on the same day, would this prove that they are exceeding their goal, using α = .05? Explain your decision. clearly state your decision on the null hypothesis.
-----------------------
Ho: p = 0.95
Ha: p > 0.95 (this implies a right-tail test)
------------------
sample proportion: (570/600) = 0.95
test statistic: z(0.95 - 0.95)/[sqrt(0.95)(0.05)/600)] = 0
--------
p-value = p(z > 0) = 0.50
---------
Since the p-value is greater than 0.05, fail to Reject Ho.
------------
Conclusion: The test does not give evidence that they are exceeding
their goal.
================================
Cheers,
Stan H.